OpenStudy (anonymous):

Which exponential function goes through the points (1, 8) and (4, 64)?

OpenStudy (anonymous):

f(x) = 4(2)^x f(x) = 2(4)^x f(x) = 4(2)^-x f(x) = 2(4)^-x